OpenStudy (coolguy34):

Please help me with this question: 3/5 + d ≤ 13/5

OpenStudy (trojanpoem):

3/5 + d <= 13/5 d <= 13- 3/5 d <= 2

OpenStudy (coolguy34):

thanks

OpenStudy (trojanpoem):

Any time.